OpenStudy (kay.baby.):

Solve x + 3 = 2x + 8

OpenStudy (kay.baby.):

How would i start?

OpenStudy (phi):

add -x to both sides

OpenStudy (tgstudios):

that's what I would have said

Awolflover1 (awolflover1):

like this \[x+3-2x=2x+8-2x\]

OpenStudy (aud101):

x+3=2x+8 x + 3 + (-x) = 2x +8 + (-x) 3+ (-8) = x+8 +(-8) -5 = x x= -5